# Reinforcement Learning for Decision-Level Interception Prioritization in Drone Swarm Defense

### v1

@@ 3.4 Environment Configuration
Sensitive zones. Zones are assumed to be circular in shape and located on the ground at $z=0.0\ m$, and specified by the location of their center $\mathbf{c}_{i}$, radius $r_{i}$, and value $v_{i}$.
@@ 4.1 Problem Formulation as Reinforcement Learning
  • Drones state, type MultiDiscrete11footnotemark: 1 with shape $[3]\times N$
  • Drones explosive, type MultiDiscrete11footnotemark: 1 with shape $[3]\times N$
@@
  • Effectors weapon state, type MultiDiscrete11footnotemark: 1 with shape $[3]\times M$
At each decision step, the agent selects one target drone per effector. The action space is therefore a MultiDiscrete11footnotemark: 1 with shape $[N]\times M$.
@@ 4.2 Preprocessing and Normalization
Box-type observations are linearly scaled to fall within the interval [1.0,0.0][-1.0,0.0][ - 1.0 , 0.0 ].
@@
Although drone impact points are initialized to be uniformly distributed across the target volume in each episode, variability in explosive payloads causes the total potential damage to vary between episodes. To address this, rewards are normalized by the theoretical maximum episode damage (i.e., if all drones reach their targets). This ensures that the total episode return remains within the range [1.0,0.0][-1.0,0.0][ - 1.0 , 0.0 ].
@@ 4.4 Training Setup and Parameters
Training is conducted using vectorized environments, with 323232 parallel simulation instances running in separate threads to accelerate data collection. The agent is trained for 808080 million steps using the Adam optimizer, with a learning rate linearly decayed from $2.5\times 10^{-4}$ to $2.5\times 10^{-6}$ and a discount factor of $\gamma=0.998$.
The PPO clipping parameter is set to 0.150.150.15 at the start of training and linearly decreased to 0.0250.0250.025. Rollout length is set to 512512512 steps, with 101010 training epochs per update and a batch size of 204820482048.
@@ 5 Evaluation & Results
$\mathcal{D}$ [100,100]×[100,100]×[0,50][-100,100$\times$[-100,100$\times$[0,50][ - 100 , 100 ] × [ - 100 , 100 ] × [ 0 , 50 ] m
$\mathcal{V}_{\text{target}}$ [100,0]×[100,100]×[0,50][-100,0$\times$[-100,100$\times$[0,50][ - 100 , 0 ] × [ - 100 , 100 ] × [ 0 , 50 ] m
$\mathcal{V}_{\text{spawn}}$ [95,100]×[100,100]×[25,50][95,100$\times$[-100,100$\times$[25,50][ 95 , 100 ] × [ - 100 , 100 ] × [ 25 , 50 ] m
$dt$ 0.10.10.1 sec
@@ 5.3 Results
Quantitatively, the RL policy achieves a 17.9617.9617.96% reduction in average damage compared to the heuristic baseline, alongside a 24.0624.0624.06% and 15.5315.5315.53% improvement in tracking efficiency and weapon utilization, respectively.
@@
Total Damage (Avg) [%] 50.34 41.30
In-Tracking Time (Avg) [%] 52.87 65.59
Weapon Utilization (Avg) [%] 54.35 62.79
Table 3: Quantitative comparison between the classical heuristic and the RL policy over 500 simulated episodes (100 per seed × 5 seeds). The RL policy significantly outperforms the heuristic across all metrics, reducing average zone damage by nearly 18%, and improving both tracking efficiency and weapon utilization.
@@ 1. Scenario
Integration Time $dt$ Constant timestep for the simulation
@@ 2. Drone Swarm
Number of Drones $N$ Total number of incoming drones in the swarm

### v2

@@ 3.4 Environment Configuration
Sensitive zones. Zones are assumed to be circular in shape and located on the ground at z=0.0mz=0.0\ m, and specified by the location of their center $\mathbf{c}_{i}$, radius $r_{i}$, and value $v_{i}$.
@@ 4.1 Problem Formulation as Reinforcement Learning
  • Drones state, type MultiDiscrete11footnotemark: 1 with shape $3]\times N$
  • Drones explosive, type MultiDiscrete11footnotemark: 1 with shape $3]\times N$
@@
  • Effectors weapon state, type MultiDiscrete11footnotemark: 1 with shape $3]\times M$
At each decision step, the agent selects one target drone per effector. The action space is therefore a MultiDiscrete11footnotemark: 1 with shape $N]\times M$.
@@ 4.2 Preprocessing and Normalization
Box-type observations are linearly scaled to fall within the interval [1.0,0.0][-1.0,0.0].
@@
Although drone impact points are initialized to be uniformly distributed across the target volume in each episode, variability in explosive payloads causes the total potential damage to vary between episodes. To address this, rewards are normalized by the theoretical maximum episode damage (i.e., if all drones reach their targets). This ensures that the total episode return remains within the range [1.0,0.0][-1.0,0.0].
@@ 4.4 Training Setup and Parameters
Training is conducted using vectorized environments, with $32$ parallel simulation instances running in separate threads to accelerate data collection. The agent is trained for $80$ million steps using the Adam optimizer, with a learning rate linearly decayed from $2.5\times 10^{-4}$ to $2.5\times 10^{-6}$ and a discount factor of $\gamma=0.998$.
The PPO clipping parameter is set to $0.15$ at the start of training and linearly decreased to $0.025$. Rollout length is set to $512$ steps, with $10$ training epochs per update and a batch size of $2048$.
@@ 5 Evaluation & Results
$\mathcal{D}$ $100,100]\times[-100,100]\times[0,50]$ m
$\mathcal{V}_{\text{target}}$ $100,0]\times[-100,100]\times[0,50]$ m
$\mathcal{V}_{\text{spawn}}$ $95,100]\times[-100,100]\times[25,50]$ m
dtdt $0.1$ sec
@@ 5.3 Results
Quantitatively, the RL policy achieves a 21.9421.94% reduction in average damage compared to the heuristic baseline, alongside a 25.3725.37% and 15.0915.09% improvement in tracking efficiency and weapon utilization, respectively.
@@
[…]
Table 3: Quantitative comparison between the classical heuristic and the RL policy over 500 simulated episodes (100 per seed × 5 seeds). The RL policy significantly outperforms the heuristic across all metrics, reducing average zone damage by nearly 22%, and improving both tracking efficiency and weapon utilization.
@@ 1. Scenario
Integration Time dtdt Constant timestep for the simulation
@@ 2. Drone Swarm
Number of Drones NN Total number of incoming drones in the swarm